### v2

# A new intersection condition in extremal set theory

Abstract.
We call a family $\mathcal{F}$ $(3,2,\ell)$-intersecting if $|A\cap B|+|B\cap C|+|C\cap A|\geq\ell$ for all $A$, $B$, $C\in\mathcal{F}$. We try to look for the maximum size of such a family $\mathcal{F}$ in case when $\mathcal{F}\subset{[n]\choose k}$ or $\mathcal{F}\subset 2^{[n]}$. In the uniform case we show that if $\mathcal{F}$ is $(3,2,2)$-intersecting, then $|\mathcal{F}|\leq{n+1\choose k-1}+{n\choose k-2}$ and if $\mathcal{F}$ is $(3,2,3)$-intersecting, then $|\mathcal{F}|\leq{n\choose k-1}+2{n\choose k-3}+3{n-1\choose k-3}$. For the lower bound we construct a $(3,2,\ell)$-intersecting family and we show that this bound is sharp when $\ell=2$ or $3$ and $n$ is sufficiently large compared to $k$. In the non-uniform case we give an upper bound for a (3,2,nx)(3,2,n-x)-intersecting family, when $n$ is sufficiently large compared to $x$.
1. Introduction
We denote the set $\{1,2,\dots,n\}$ by $[n]$, the family of the subsets of $[n]$ by $2^{[n]}$, and the family of the $i$-element subsets of $[n]$ by $\binom{[n]}{i}$. If $i>n$, then this notation means the empty set.
We will mostly use the terminology families for collections of sets, although set systems is also widely used in the literature. If all sets in a family $\mathcal{F}$} are of the same size, we say $\mathcal{F}$} is uniform. Otherwise $\mathcal{F}$} is called non-uniform.
Definition 1.1.
A family $\mathcal{F}$} is said to be intersecting if $|A\cap B|\neq 0$ holds for all $A$, $B\in{\mathcal{F}}$.
The following two theorems of Erdős, Ko and Rado are central ones in extremal set theory. They determine the largest size of a family containing no disjoint pairs of sets. The first theorem solves the case of non-unifom families, the second one determines the maximum for the uniform case.
Theorem 1.2.
[2] Let ${\mathcal{F}}\subset 2^{[n]}$ be an intersecting family. Then $|{\mathcal{F}}|\leq 2^{n-1}$.
Theorem 1.3.
[2] Let $k$, $n\geq 2k$ be integers and ${\mathcal{F}}\subset{[n]\choose k}$ be an intersecting family. Then $|{\mathcal{F}}|\leq\displaystyle\binom{n-1}{k-1}$.
The concept "intersecting family" has several generalizations, as follows.
Definition 1.4.
A family $\mathcal{F}$ is said to be $t$-intersecting if $|A\cap B|\geq t$ holds for all $A$, $B\in\mathcal{F}$.
Definition 1.5.
A family $\mathcal{F}$ is said to be trivially $t$-intersecting if there exists a $t$-set $T$ such that $T\subset F$ holds for all $F\in\mathcal{F}$, and non-trivially $t$-intersecting if it is $t$-intersecting but $|\bigcap_{F\in\mathcal{F}}F|<t$.
It was proved by Erdős, Ko, and Rado in their seminal paper [2] that for any $1\leq t<k$, the largest $t$-intersecting families $\mathcal{F}\subset{[n]\choose k}$ are the maximal trivially $t$-intersecting ones if $n$ is large enough. The threshold function $n_{0}(k,t)=(k-t+1)(t+1)$ of this property was determined by Frankl [5] for $t\geq 15$. Later Wilson [19] filled in the gap $2\leq t\leq 14$.
Theorem 1.6.
Trivially $t$-intersecting family is the maximal $t$-intersecting family if and only if $n\geq(t+1)(k-t+1)$.
The non-uniform version of the theorem for $t\geq 2$ was proved by Katona.
Theorem 1.7.
[16] Let ${\mathcal{F}}\subset 2^{[n]}$ be a $t$-intersecting family with $1\leq t\leq n$. Then
$|\mathcal{F}|\leq\left\{\begin{array}[]{c c}\sum_{i=\frac{n+t}{2}}^{n}{n\choose i}&\text{if $n+t$ is even}\vskip 2.84526pt\\ \vskip 2.84526pt\sum_{i=\frac{n+t+1}{2}}^{n}{n\choose i}+{n-1\choose\frac{n+t-1}{2}}&\text{if $n+t$ is odd}\end{array}\right$.
One of the famous questions of this area is the Erdős Matching Conjecture (EMC), in which the maximum number of $k$-element subsets should be determined under the condition that no $s+1$ of them are pairwise disjoint. This value is denoted by $m(n,k,s)$.
Conjecture 1.8.
[3] For $n\geq k(s+1)$,
$m(n,k,s)=\max\left\{{n\choose k}-{n-s\choose k},{k(s+1)-1\choose k}\right\}$.
It was one of the favourite problems of Erdős and there was hardly a combinatorial lecture of him where he did not mention it. The EMC is trivial for $k=1$ and was proved by Erdős and Gallai [4] for $k=2$ then it was settled in the case $k=3$ and $n\geq 4s$ in [13], for $k=3$, all $n$ and $s\geq s_{0}$ in [18], and finally, it was completely resolved for $k=3$ in [8].
The case $s=1$ is the classical Erdős-Ko-Rado theorem. In his original paper, Erdős proved 1.8 for $n\geq n_{0}(k,s)$ for some $n_{0}(k,s)$. His result was sharpened by Bollobás, Daykin and Erdős [1], they proved for $n\geq 2k^{3}s$. Subsequently, Hao, Loh and Sudakov [15] proved the EMC for $n\geq 3k^{2}s$.
Frankl [9] showed that $m(n,k,s)\leq{n\choose k}-{n-s\choose k}$ when $n\geq(2s+1)k-s$. For $s$ very large Frankl and Kupavskii prove the following theorem.
Theorem 1.9.
[11] There exists an absolute constant $s_{0}$, such that
$m(n,k,s)={n\choose k}-{n-s\choose k}$
holds if $n\geq\frac{5}{3}sk-\frac{2}{2}s$ and $s\geq s_{0}$.
In the non-uniform case Kleitman considered when $s=2$.
Theorem 1.10.
[17] Let $n=3m+2$ and $\mathcal{F}\subset 2^{[n]}$ be a family with no three of whose members are pairwise disjoint. Then
$|\mathcal{F}|\leq\sum_{j=m+1}^{3m+2}{3m+2\choose j}$
is the best upper bound.
Theorem 1.11.
[17] Let $n=3m$ and $\mathcal{F}\subset 2^{[n]}$ be a family with no three of whose members are pairwise disjoint. Then
$|\mathcal{F}|\leq\sum_{j=m+1}^{3m}{3m\choose j}+\frac{2}{3}{3m\choose m}$
is the best upper bound.
2. Results
The condition of Kleitman’s theorem can be written in the following form: $|A\cap B|+|B\cap C|+|C\cap A|\geq 1$ holds for any 3 distinct members $A$, $B$, $C\in{\mathcal{F}}$. The main goal of the present paper is to study the case when 1 is replaced by $\ell$. Since the case $\ell=1$ has been studied a lot, we will only look at the case $\ell\geq 2$.
Let $d:(2^{[n]})^{3}\rightarrow\mathbb{N}$ be a function such that for all $A$, $B$, $C\in 2^{[n]}$, $d(A,B,C)=|A\cap B|+|B\cap C|+|C\cap A|$. Denote by $\mathbf{H}_{\ell}$ the class of families such that $\mathcal{H}$, $d(A,B,C)\geq\ell$ for all sets $A$, $B$, $C\in\mathcal{H}$. Let $g(n,k,\ell)=\max\{|\mathcal{F}|:\mathcal{F}\subset{[n]\choose k},\mathcal{F}\in\mathbf{H_{\ell}}\}$ and $h(n,\ell)=\max\{|\mathcal{F}|:\mathcal{F}\subset 2^{[n]},\mathcal{F}\in\mathbf{H_{\ell}}\}$.
First we consider the uniform case. We prove the following upper and lower bounds:
Theorem 2.1.
Let $n\geq 3k-2\geq 4$ be positive integers. Then
$g(n,k,2)\leq{n+1\choose k-1}+{n\choose k-2}$.
Theorem 2.2.
Let $n\geq 3k-3\geq 3$ be positive integers. Then
$g(n,k,3)\leq{n\choose k-1}+2{n\choose k-3}+3{n-1\choose k-3}$.
Theorem 2.3.
Let $k$, $\ell$, $n$ be positive integers with $3k\geq\ell$. Then
(1) $g(n,k,\ell)\geq\max_{k\geq j\geq\frac{\ell}{3}}\left\{\sum_{i=j}^{k}{f(j,\ell)\choose i}{n-f(j,\ell)\choose k-i}\right\}$,
where
[…]
Remark 2.4.
If $a<b$, then we interpret $a\choose b$} as zero.
Finally, we show that if $\ell=2$ or $3$ and $n$ is sufficiently large compared to $k$, then our upper bounds in Theorem 2.1 and Theorem 2.2 can be improved, reaching the value of the lower bound in Theorem 2.3 for $j=1$.
Theorem 2.5.
Let $n\geq 4\cdot k^{3}$ be positive integers. Then $g(n,k,2)=g(n,k,3)={n-1\choose k-1}$.
This theorem is a sharpening of the Erdős-Ko-Rado theorem, since it achieves the same upper bound for the corresponding pairs $(n,k)$ under weaker conditions.
In the non-uniform case, we investigate families $\mathcal{F}$ for which $d(A,B,C)\geq 3n-x$ for all $A$, $B$, $C\in\mathcal{F}$. In the case when $x=6p$ ($p$ is an integer) the following construction seems to be optimal: we select all sets with at least npn-p elements, then for any set of triples $d(A,B,C)\geq 3n-6p$. If $x=6p$, this will be the optimal solution for large $n$ and if $6p<x<6(p+1)$, then the optimal solution contains all npn-p-element sets and some of the np1n-p-1-element sets.
Theorem 2.6.
Let $x=6p+q$, where $1\leq p$, $0\leq q\leq 5$ and $n\geq 2^{3p+2}p^{2}+p+1$. Then
[…]
3. The uniform case
In the next section, we use Frankl’s shifting technique. [10]
3.1. Shifting a family
Notation.
Let $\mathcal{A}$ be a three-member family $\mathcal{A}=\{A,B,C\}$.
Definition 3.1.
Let $1\leq y<x\leq n$ be two elements of $[n]$ and let $\mathcal{F}\subset 2^{[n]}$ be a family. Then
$\mathcal{F}^{\prime}=\{F:F\in\mathcal{F},(x\notin F)\vee(x,y\in F)\vee(x\in F,y\notin F,F\cup\{y\}-\{x\}\in\mathcal{F})\}$
and
$\mathcal{F}^{\prime\prime}=\{F:F\in\mathcal{F},x\in F,y\notin F,F\cup\{y\}-\{x\}\notin\mathcal{F}\}.$
Definition 3.2.
Let $x$, $y$ and $\mathcal{F}$ be same as above. Then
$\tau_{x,y}(\mathcal{F})=\mathcal{F}^{\prime}\cup\{F\cup\{y\}-\{x\}:F\in\mathcal{F}^{\prime\prime}\}.$
This operation is called a shifting of $\mathcal{F}$ (from $x$ to $y$).
Definition 3.3.
Let $\tau_{x,y}:\mathcal{F}\rightarrow\tau_{x,y}(\mathcal{F})$ be a bijection such that for $\forall F\in\mathcal{F}$
(2) $\tau_{x,y}(F)=\begin{cases}F&\text{if $(x,y\in F)\vee(x,y\notin F)$}\\ F&\text{if $x\in F,y\notin F,F\cup\{y\}-\{x\}\in\mathcal{F}$ }\\ F-\{x\}\cup\{y\}&\text{otherwise}\end{cases}$
Claim 3.4.
Let $\mathcal{F}\in 2^{[n]}$ be a family such that $d(\mathcal{A})\geq\ell$ for all $A$, $B$, $C\in\mathcal{F}$. Then $d(\tau_{x,y}(A),\tau_{x,y}(B),\tau_{x,y}(C))\geq\ell$.
Notation.
When it does not cause ambiguity we delete the indices $x$, $y$ from the notation $\tau_{x,y}$.
Proof.
We prove this claim by contradiction. Suppose that the conditions are true, but there exist sets $A$, $B$ and $C\in{\mathcal{F}}$ such that $d(\tau(\mathcal{A}))<\ell$.
Suppose that $z\notin\{x,y\}$. Then $z\in F$ holds if and only if $z\in\tau(F)$ so $d(\mathcal{A}\setminus\{x,y\})=d(\tau(\mathcal{A})\setminus\{x,y\})=c$.
Let ${\mathcal{F}}^{1}=\{H:H\in{\mathcal{F}},H=\tau(H)\}$ and ${\mathcal{F}}^{2}=\{H:H\in{\mathcal{F}},H\neq\tau(H)\}$. Let $\mathcal{A}^{1}=\mathcal{A}\cap{\mathcal{F}}^{1}$ and $\mathcal{A}^{2}=\mathcal{A}\cap{\mathcal{F}}^{2}$. We can assume that $\mathcal{A}^{1}$ contains the element $x$ $p$ times, the element $y$ $q$ times and $\mathcal{A}^{2}$ contains the element $x$ $r$ times. (It is obvious that $\mathcal{A}^{2}$ contain the element $y$ 0 times.) Then $d(\mathcal{A})=c+{p+r\choose 2}+{q\choose 2}>c+{p\choose 2}+{q+r\choose 2}=d(\tau(\mathcal{A}))$.
This is true only if $p+r>p>q$ (and $p+r>q+r>q$). This means that there are at least pqp-q sets $H\in\mathcal{A}^{1}$ such that $x\in H$, $y\notin H$. Since $x\in\tau(H)$ and $y\notin\tau(H)$, then $H^{\prime}=H-\{x\}\cup\{y\}\in{\mathcal{F}}$. So if $\mathcal{A}^{\prime}$ is a family such that we replace $x$ with $y$ in this pqp-q sets in $\mathcal{A}^{1}$, then $\mathcal{A}^{\prime}\subset{\mathcal{F}}$ and $d(\mathcal{A}^{\prime})=c+{p\choose 2}+{q+r\choose 2}=d(\tau(\mathcal{A}))<\ell$, a contradiction. ∎
Definition 3.5.
Let $H\in 2^{[n]}$ and $\mathcal{H}\subset 2^{[n]}$, then $s(H)=\sum_{i\in H}i$ and $s(\mathcal{H})=\sum_{H\in\mathcal{H}}s(H).$
Claim 3.6.
If $\tau(\mathcal{H})\neq\mathcal{H}$, then $s(\tau(\mathcal{H}))<s(\mathcal{H})$. $\square$
Definition 3.7.
Let $H=\{h_{1},\dots,h_{k}\}$ and $I=\{i_{1}\dots i_{k}\}\in{[n]\choose k}$ be sets such that $h_{1}<h_{2}<\dots<h_{k}$ and $i_{1}<i_{2}<\dots<i_{k}$. We write that $I\leq H$, if $i_{j}\leq h_{j}$ holds for all $1\leq j\leq k$.
Definition 3.8.
A uniform family $\mathcal{F}\in{[n]\choose k}$ is called shifted, if $H\in\mathcal{F}$ and $I\leq H$ imply $I\in\mathcal{F}$. A non-uniform family $\mathcal{F}\in 2^{[n]}$ is called shifted, if $\mathcal{F}\cap{[n]\choose i}$ is shifted for every $0\leq i\leq n$.
Claim 3.9.
For a family $\mathcal{F}\in 2^{[n]}$, there exist $x$ and $y$ such that $s(\mathcal{F})>s(\tau_{x,y}(\mathcal{F}))$ if and only if $\mathcal{F}$ is not shifted. $\square$
Claim 3.10.
For a family $\mathcal{F}\in 2^{[n]}$, there exist $\tau_{1},\dots,\tau_{k}$ such that $\mathcal{F^{\prime}}=\tau_{x_{1},y_{1}}(\tau_{x_{2},y_{2}}(\dots(\tau_{x_{k},y_{k}}(\mathcal{F}))))$ is a shifted family.
Proof.
If $\mathcal{F}$ is not shifted, there exists a $\tau$ shifting such that $s(\mathcal{F})>s(\tau(\mathcal{F}))$. Since $s(\mathcal{H})\in\mathbb{N}$, by Claim 3.6 there cannot be an infinite decreasing sequence, so after a finite number of shifting we obtain a shifted family. ∎
Corollary 3.11.
As a result of Claim 3.4 and 3.10, if a family $\mathcal{F}\in\mathbf{H}_{\ell}$, then there exists a shifted family $\mathcal{F^{\prime}}\in\mathbf{H}_{\ell}$ such that $|\mathcal{F}|=|\mathcal{F}^{\prime}|$. Therefore it is sufficient to consider the shifted families, when we search for the maximum sized family in $\mathbf{H}_{\ell}$.
Claim 3.12.
Let $\ell=2$ or $3$, $\mathcal{F}\subset{[n]\choose k}$ and $\mathcal{F}\in\mathbf{H}_{\ell}$ be a shifted family. If $A$, $B$, $C\in\mathcal{F}$ and $A^{\prime}=A\cap[3k-\ell]$, $B^{\prime}=B\cap[3k-\ell]$ and $C^{\prime}=C\cap[3k-\ell]$, then $d(A^{\prime},B^{\prime},C^{\prime})\geq\ell$, that is $\mathcal{F}\cap[3k-\ell]\in\mathcal{H}_{\ell}$.
Proof.
Let $A_{1}=A\setminus[3k-\ell]$, $B_{1}=B\setminus[3k-\ell]$ and $C_{1}=C\setminus[3k-\ell]$.
Suppose that the statement is not true and let $A$, $B$, $C\in\mathcal{F}$ such that $d(A^{\prime},B^{\prime},C^{\prime})<\ell$ and $m=|A_{1}|+|B_{1}|+|C_{1}|$ is the smallest possible. Since $d(A_{1},B_{1},C_{1})>0$, so $m\geq 2$ and wlog $p\in A_{1}$.
If there is an $a\in[3k-\ell]\setminus(A^{\prime}\cup B^{\prime}\cup C^{\prime})$ then $\overline{A}=A-\{p\}\cup\{a\}\in\mathcal{F}$, $d(\overline{A}^{\prime},B^{\prime},C^{\prime})=d(A^{\prime},B^{\prime},C^{\prime})<\ell$ and $|\overline{A}_{1}|+|B_{1}|+|C_{1}|<m$ which is impossible.
So $A^{\prime}\cup B^{\prime}\cup C^{\prime}=[3k-\ell]$. Since $d(A^{\prime},B^{\prime},C^{\prime})<\ell\leq 3$, then there is no $a\in[3k-\ell]$ such that $a\in A\cap B\cap C$. Since $|A^{\prime}|+|B^{\prime}|+|C^{\prime}|=3k-m\leq 3k-2$, then $d(A^{\prime},B^{\prime},C^{\prime})=\ell-m\leq\ell-2$. Let $q\in[3k-\ell]$ so that $q$ is in one of $A^{\prime}$, $B^{\prime}$, $C^{\prime}$. Then $\overline{A}=A-\{p\}\cup\{q\}\in F$ and $d(\overline{A^{\prime}},B^{\prime},C^{\prime})\leq\ell-2+1<\ell$ but $|\overline{A}_{1}|+|B_{1}|+|C_{1}|<m$ which is a contradiction. ∎
Remark 3.13.
Claim 3.12 is not true if $\ell\geq 4$. For example let $\ell=4$, $k=3$, $A=\{1,2,6\}$, $B=\{1,3,6\}$, $C=\{1,4,5\}$ and $H\in\mathcal{F}$ if and only if $H\leq A$ or $H\leq B$ or $H\leq C$. Then $\mathcal{F}\in\mathbf{H}_{\ell}$, because for all $H\in{\mathcal{F}}$ are true that $1\in H$ and contains two element of $\{2,3,4,5,6\}$, so $d(\mathcal{A}\cap{1})=3$ and $d(\mathcal{A}-{1})\geq 1$, so $d(\mathcal{A})\geq 4$ but $3k-\ell=5$ and $d(A^{\prime},B^{\prime},C^{\prime})<4$.
Claim 3.14.
If $\ell=2$ or $\ell=3$, $k\geq 3$ and $n>3k$-$\ell$, then $g(n,k,\ell)\leq g(n-1,k,\ell)+g(n-1,k-1,\ell)$.
Proof.
Let $\mathcal{M}_{n,k,\ell}\subset{[n]\choose k}$ be a shifted family such that $\mathcal{M}_{n,k,\ell}\in\mathbf{H}_{\ell}$ and $|\mathcal{M}_{n,k,\ell}|=g(n,k,\ell)$. This exists because of Corollary 3.11.
Let $\mathcal{M}^{1}_{n,k,\ell}\subset{[n-1]\choose k}$ be the family $\mathcal{M}^{1}_{n,k,\ell}=\{M:M\in\mathcal{M}_{n,k,\ell},\{n\}\notin M\}$ and let $\mathcal{M}^{2}_{n,k,\ell}\subset{[n-1]\choose k-1}$ be the family such that $\mathcal{M}^{2}_{n,k,\ell}=\{M\setminus\{n\}:M\in\mathcal{M}_{n,k,\ell},n\in M\}$.
It is easy to check that $\mathcal{M}^{1}_{n,k,\ell}$ and $\mathcal{M}^{2}_{n,k,\ell}$ are shifted families. Moreover $\mathcal{M}^{1}_{n,k,\ell}\subset\mathcal{M}_{n,k,\ell}$, so $\mathcal{M}^{1}_{n,k,\ell}\in\mathbf{H}_{\ell}$. Since $n>3k$-$\ell$, $\mathcal{M}^{2}_{n,k,\ell}\cap[3k-\ell]\subset\mathcal{M}_{n,k,\ell}\cap[3k-\ell]$. Using Claim 3.12, we obtain that $\mathcal{M}_{n,k,\ell}\in\mathbf{H}_{\ell}$ implies $\mathcal{M}^{2}_{n,k,\ell}\in\mathbf{H}_{\ell}$.
Since $|\mathcal{M}^{1}_{n,k,\ell}|\leq g(n-1,k,\ell)$ and $|\mathcal{M}^{2}_{n,k,\ell}|\leq g(n-1,k-1,\ell)$, so $g(n,k,\ell)=|\mathcal{M}_{n,k,\ell}|=|\mathcal{M}^{1}_{n,k,\ell}|+|\mathcal{M}^{2}_{n,k,\ell}|\leq g(n-1,k,\ell)+g(n-1,k-1,\ell)$. ∎
3.2. The proof of Theorem 2.1 and 2.2
Proof.
It is easy to see that $g(3k-\ell,k,\ell)={3k-\ell\choose k}$. In the next step we show that Theorem 2.1 and Theorem 2.2 are true when $k=2$.
Claim 3.15.
$g(n,2,2)\leq{n+1\choose 2-1}+{n\choose 2-2}=n+2$.
Proof.
The statement is trivially true when $n\leq 3$, so we assume that $n\geq 4$. If there are $A$ and $B$ in $\mathcal{F}$ such that $|A\cap B|=0$, then for all $C\in\mathcal{F}$, $C\subset A\cup B$. Since $|A\cup B|=4$, then $|\mathcal{F}|\leq 6\leq n+2$. Therefore, we can assume that such $A$ and $B$ do not exist. So $A=\{a,b\}$ and $B=\{a,c\}$. Then, for all $C\in\mathcal{F}$, either $a\in C$ or ${b,c}\in C$. Therefore, $|\mathcal{F}|\leq n-1+1=n$. ∎
Claim 3.16.
$g(n,2,3)\leq{n\choose 2-1}=n$.
Proof.
Here we can assume that $n\geq 3$. If there exist $A$, $B\in\mathcal{F}$ such that $|A\cap B|=0$, then for all $C\in\mathcal{F}$, $d(A,B,C)\leq 2$, so $|\mathcal{F}|\leq 2$. If $A\cap B\neq\emptyset$ for all $A$ and $B$ in $\mathcal{F}$, then fix $A=\{a,b\}$ and $B=\{a,c\}$. Then for all $C\in\mathcal{F}$, either $a\in C$ or ${b,c}\in C$, so $|\mathcal{F}|\leq n-1+1=n$. ∎
From here we prove both theorems by induction on $n$ and $k$. We have already proven the statement if either $k=2$ or $n=3k$-$\ell$. Now, let us consider the case where $k\geq 3$ and $n>3k$-$\ell$.
According to Claim 3.14 we know that $g(n,k,2)\leq g(n-1,k,2)+g(n-1,k-1,2)={n\choose k-1}+{n-1\choose k-2}+{n\choose k-2}+{n-1\choose k-3}={n+1\choose k-1}+{n\choose k-2}$ and $g(n,k,3)\leq g(n-1,k,3)+g(n-1,k-1,3)={n-1\choose k-1}+2{n-1\choose k-3}+3{n-2\choose k-3}+{n-1\choose k-2}+2{n-1\choose k-4}+3{n-2\choose k-4}={n\choose k-1}+2{n\choose k-3}+3{n-1\choose k-3}$, so the theorems are true. ∎
3.3. Lower bound for the uniform case
Definition 3.17.
Let $f(\mathcal{H},m)$ denote the number of elements of $[n]$ contained in exactly $m$ members of $\mathcal{H}$.
This definition leads to the following easily verifiable statement:
Claim 3.18.
For all $\mathcal{A}\in\left({2^{[n]}}\right)^{3}$ we have $d(\mathcal{A})=3\cdot f(\mathcal{A},3)+f(\mathcal{A},2)$.
Using this, we can give a lower bound on $d(\mathcal{A})$.
Claim 3.19.
Let $\mathcal{A}\in\left({2^{[n]}}\right)^{3}$ be a family and let the sum of their sizes be $s=|A|+|B|+|C|$. Now if $n\leq h\leq 2n$ then
(3) $d(\mathcal{A})\geq s-n$
if $2n\leq s\leq 3n$ then
(4) $d(\mathcal{A})\geq n+2\cdot(s-2n)$
hold.
Proof.
It is obvious that
(5) $s=3f(\mathcal{A},3)+2f(\mathcal{A},2)+f(\mathcal{A},1)$
and hence we have
(6) $d(\mathcal{A})=3f(\mathcal{A},3)+f(\mathcal{A},2)=s-f(\mathcal{A},2)-f(\mathcal{A},1).$
Here $f(\mathcal{A},2)+f(\mathcal{A},1)\leq n$ is obvious and (6) leads to (3).
In order to prove (4) let us start with the inequality $3n\geq 2f(\mathcal{A},1)+3f(\mathcal{A},2)+3f(\mathcal{A},3)$. Using (5) the right-hand side can be written in the following form: $3n\geq f(\mathcal{A},1)+f(\mathcal{A},2)+s$.
This is equivalent to the inequality
(7) $s-f(\mathcal{A},1)-f(\mathcal{A},2)\geq n+2(s-2n).$
Here the left hand side is just $d(\mathcal{A})$ by (6) and (7) really gives (4). ∎
Claim 3.20.
Let $j$ and $\ell\leq 3j$ be integers and let $A$, $B$, $C\in\displaystyle\binom{[f(j,\ell)]}{j}$ be three sets. Then $d(\mathcal{A})\geq\ell.$
Proof.
Case 1: $\frac{\ell}{3}\leq j<\frac{2\ell}{3}$. Using the notation of 3.19, $s=3j$ and $n=f(j,l)=2j-\lceil\frac{\ell}{3}\rceil$. Since $2(2j-\lceil\frac{\ell}{3}\rceil)\leq 3j\leq 3(2j-\lceil\frac{\ell}{3}\rceil)$ we can use (4), so $d(\mathcal{A})\geq 2s-3n=2\cdot 3j-3(2j-\lceil\frac{\ell}{3}\rceil)\geq\ell$.
Case 2: $\frac{2\ell}{3}\leq j$. Using the same notation, $h=3j$ and $n=f(j,\ell)=3j$-$\ell$. Since $3j-\ell\leq 3j\leq 2(3j-\ell)$ we can use (3), so $d(\mathcal{A})\geq s-n=3j-(3j-\ell)=\ell$. ∎
Proof of Theorem 2.3. Let $j$ be an integer and let $\mathcal{F}=\{F:F\in{[n]\choose k},|F\cap[f(j,\ell)]|\geq j\}$ be a family. Because of Claim 3.20 we know that $\mathcal{F}\in\mathbf{H}_{\ell}$. The expression ${f(j,\ell)\choose i}{n-f(j,\ell)\choose k-i}$ is the number of sets containing $i$ elements from $3j-\ell]$. By choosing the optimal $j$ we get the lower bound. $\square$
Remark 3.21.
If $3k\leq 2n$ and $\ell\leq 3k-n$ or $2n<3k$ and $\ell\leq 9k-5n$, then $g(n,k,\ell)={n\choose k}$.
Conjecture 3.22.
In cases that are not included in the Remark 3.21, the lower bound in Theorem 2.3 is sharp.
For a given $(n,k,\ell)$, it is difficult to compute the value (1) in Theorem 2.3, as it must be determined for each $j$. In the rest of the section, we make observations about the expected optimal value of $j$.
Conjecture 3.23.
For every $\ell\geq 2$ there is a sequence $0=\alpha_{0}(\ell)<\alpha_{1}(\ell)<\alpha_{2}(\ell)<\dots<\frac{1}{3}$ of real numbers such that if $n$, $k$ tends to infinity so that $\frac{k}{n}\rightarrow\gamma$ where $\alpha_{i-1}(\ell)<\gamma<\alpha_{i}(\ell)$ then the right-hand side of (1) gives the maximum for $j=i$.
Definition 3.24.
Let $m(n,k,\ell)$ be the minimum value of $j$ where (1) gives the maximum.
In the following claim we determine the value of $\alpha_{1}(2)$ in Conjecture 3.23.
Claim 3.25.
Let $\ell=2$ and $n$ be sufficiently large. Then $m(n,k,\ell)\geq 2$ if the ratio $\frac{k}{n}$ is greater then $\frac{5-\sqrt{13}}{6}$.
Proof.
If $j=1$, then $f(j,\ell)=1$, so the value of the expression in (1) is $n-1\choose k-1$. If $j=2$, then $f(j,\ell)=4$, so in this case the value of the expression in (1) is ${4\choose 2}{n-4\choose k-2}+{4\choose 3}{n-4\choose k-3}+{4\choose 4}{n-4\choose k-4}$. So $m(n,k,\ell)\geq 2$ if
$\frac{(n-1)(n-2)(n-3)}{(k-3)(k-2)(k-1)(n-k-1)(n-k)}\leq 6\cdot\frac{1}{(k-3)(k-2)}+$
$4\cdot\frac{1}{(k-3)(n-k-1)}+\frac{1}{(n-k-1)(n-k)}.$
If we take $n$ (and $k$) to infinity, we get this approximation
$\frac{n^{3}}{k^{3}(n-k)^{2}}\leq 6\cdot\frac{1}{k^{2}}+4\cdot\frac{1}{k(n-k)}+\frac{1}{(n-k)^{2}}.$
Multiplying by the denominators it gives $n^{3}\leq 6\cdot k\cdot n^{2}-8\cdot k^{2}\cdot n+3\cdot k^{3}$. This means that $3\left(\frac{k}{n}\right)^{3}-8\left(\frac{k}{n}\right)^{2}+6\left(\frac{k}{n}\right)-1\geq 0$. It is true, if $1\geq\frac{k}{n}\geq\frac{5-\sqrt{13}}{6}$ or $0\geq\frac{k}{n}$, so $m(n,k,\ell)\geq 2$ if $\frac{k}{n}\geq\frac{5-\sqrt{13}}{6}\pm\varepsilon\approx 0.23$. ∎
In a similar way we get that $m(n,k,\ell)\geq 3$ if and only if $\frac{k}{n}\geq\alpha_{2}\approx 0.29$.
Proof of Theorem 2.5. First let us assume that there are two sets $A$ and $B\in\mathcal{F}$ that are disjoint. Since $|A\cap B|=0$, therefore $|A\cap C|+|B\cap C|\geq 2$, so all sets must contain at least 2 elements from $A\cup B$. An upper bound can be obtained by choosing the first two elements of the sets from $A\cup B$ and the others from $[n]$, so that $|\mathcal{F}|\leq{2k\choose 2}{n\choose k-2}$.
The other case when there are no two disjoint sets. In that case we can use Theorem 1.3, so $|\mathcal{F}|\leq{n-1\choose k-1}$ and equality is achieved if and only if there is an element $p\in[n]$ such that $H\in\mathcal{F}$ if and only if $p\in H$.
To prove the theorem we have to show that ${n-1\choose k-1}\geq{2k\choose 2}{n\choose k-2}$. Simplifying this, we get that $(n-k+2)(n-k+1)\geq nk(k-1)(2k-1)$. Using the condition $n\geq 4k^{3}$, we get that $(n-k+2)(n-k+1)\geq\frac{n^{2}}{2}\geq\frac{n}{2}4k^{3}\geq nk(k-1)(2k-1)$. $\square$
4. The non-uniform case
4.1. General observations in the non-uniform case
Definition 4.1.
Let $\mathcal{F}\subset 2^{n}$, $\subset G, FF\in\mathcal{F}$ and $G\notin\mathcal{F}$. We call $\mathcal{F}-\{F\}\cup\{G\}$ as an upward-shifting of $\mathcal{F}$. A family $\mathcal{F^{\prime}}$ is called an up-shifted version of $\mathcal{F}$ if $\mathcal{F^{\prime}}$ can be obtained from $\mathcal{F}$ by some upward-shifting.
Definition 4.2.
We call a family $\mathcal{F}\subset 2^{[n]}$ upward-closed if for every $F\in\mathcal{F}$, if $F\subset G$, then $G\in\mathcal{F}$.
The following two claims are easy to prove.
Claim 4.3.
For each family $\mathcal{F}\subset 2^{[n]}$, there exists an upward-closed family $\mathcal{F^{\prime}}\subset 2^{[n]}$ which is an up-shifted version of $\mathcal{F}$. $\square$
Claim 4.4.
Let $\mathcal{F}^{\prime}$ be an up-shifted version of $\mathcal{F}$. Then if $d(\mathcal{A})\geq\ell$ for all $\mathcal{A}\subset\mathcal{F}$, then $d(A^{\prime},B^{\prime},C^{\prime})\geq\ell$ for all $\{A^{\prime},B^{\prime},C^{\prime}\}=\mathcal{A^{\prime}}\subset\mathcal{F}^{\prime}$. $\square$
Claim 4.5.
Suppose that the family $\mathcal{F}$ is such that $d(\mathcal{A})\geq\ell$ is true for all set-triples $\mathcal{A}\subset\mathcal{F}$ then there exists an upward-closed shifted family $\mathcal{F^{\prime}}$ such that for all sets $\mathcal{A^{\prime}}\subset\mathcal{F^{\prime}}$, $d(\mathcal{A^{\prime}})\geq\ell$ and $|\mathcal{F}|=|\mathcal{F^{\prime}}|$ are satisfied.
Proof.
Take an upward-closed family $\mathcal{F}_{1}$ which is up-shifted of $\mathcal{F}$. It exists because of Claim 4.3. If it is a shifted family, then we are done. If not, then we can take a shifted family $\mathcal{F}_{2}$ which is a shift of $\mathcal{F}_{1}$ because of Claim 3.10. If it is an upward-closed family, then we are done. If not, we can do this cycle again.
We can see that $\sum_{F\in\mathcal{F}}|F|$ increases with each upward-shifting and remains the same with each shift. Since $\sum_{F\in\mathcal{F}}|F|$ has an upper bound, so the cycle once it stops and then we get an upward-closed shifted family for which the condition is true because of Corollary 3.11 and Claim 4.4. ∎
4.2. Sets with large intersection
Notation.
Let $\overline{d(\mathcal{A})}=2f(\mathcal{A},1)+3f(\mathcal{A},2)+3f(\mathcal{A},3)$ and $\overline{d(A,B)}=2f(\{A,B\},1)+3f(\{A,B\},2)$. Denote by $\overline{\mathbf{H}_{\ell}}$ the class of families such that $\mathcal{H}\in\overline{\mathbf{H}_{\ell}}$ and $\mathcal{A}\subset\mathcal{H}$ imply $\overline{d(\mathcal{A})}\leq\ell$. Let $\mathcal{F}^{-}=\{\overline{F}:F\in\mathcal{F}\}$ and let $\overline{h}(n,\ell)=\max\{|\mathcal{F}|:\mathcal{F}\subset 2^{[n]},\mathcal{F}\in\overline{\mathbf{H}_{\ell}}\}$.
Remark 4.6.
The following proof is not working when $p=0$. It is easy to see that in this case h(n,3nx)=2h(n,3n-x)=2 if $0\leq x\leq 3$ and h(n,3nx)=3h(n,3n-x)=3 if $4\leq x\leq 5$.
Proof of Theorem 2.6. The following claims are needed to prove the theorem.
Claim 4.7.
Let $A$, $B$, $C$ be sets such that $|A|+|B|+|C|=s$. Then $d(\mathcal{A})+\overline{d(\mathcal{A})}=2s$.
Proof.
$d(\mathcal{A})+\overline{d(\mathcal{A})}=2f(\mathcal{A},1)+4f(\mathcal{A},2)+6f(\mathcal{A},3)=2s$. ∎
Claim 4.8.
$h(n,3n-\ell)=max\{|\mathcal{F}|:\mathcal{F}\subset 2^{[n]},\mathcal{F}\in\overline{\mathbf{H_{\ell}}}\}=\overline{h}(n,\ell)$.
Proof.
From Claim 3.18 we know that $d(\mathcal{A})=3f(\mathcal{A},3)+f(\mathcal{A},2)$ and $f(\mathcal{A},3)+f(\mathcal{A},2)+f(\mathcal{A},1)+f(\mathcal{A},0)=n$, so $d(\mathcal{A})=3n-2f(\mathcal{A},2)-3f(\mathcal{A},1)-3f(\mathcal{A},0)=3n-2f(\mathcal{A}^{-},1)-3f(\mathcal{A}^{-},2)-3f(\mathcal{A}^{-},3)=3n-\overline{d(\mathcal{A}^{-})}$.
We obtained that $d(\mathcal{A})+\overline{d(\mathcal{A}^{-})}=3n$ holds for all $\mathcal{A}\subset\mathcal{F}$ consequently $\mathcal{F}\in\mathbf{H_{\ell}}$ if and only if $\mathcal{F}^{-}\in\overline{\mathbf{H_{\ell}}}$. ∎
Using this claim, we obtain the following modified problem. Maximize $|\mathcal{F}|$ under the condition that all $\mathcal{A}\subset\mathcal{F}$ satisfy $\overline{d(\mathcal{A})}\leq x$.
We will distinguish cases according to the divisibility of $x$ by 6. Let $x=6p+q$, where $p$, $q$ are integers and $0\leq q\leq 5$.
Claim 4.9.
Let $x\geq 6p$. Then $\overline{h}(n,x)\geq\sum_{i=0}^{p}{n\choose i}$.
Proof.
Let $\mathcal{F}=\{F:|F|\leq p\}$ be a family. By Claim 4.7, for all $\mathcal{A}\subset\mathcal{F}$, $\overline{d(\mathcal{A})}\leq 2(|A|+|B|+|C|)=6p$, which means $\mathcal{F}\in\overline{\mathbf{H}_{6p}}\subset\overline{\mathbf{H}_{x}}$. ∎
Claim 4.10.
Let $x=6p+q$ and $n\geq 2^{3p+2}p^{2}+p+1$. If there exists $A$, $B\in\mathcal{F}$ such that $\overline{d(A,B)}>4p+q$, then $\mathcal{F}$ cannot be a maximal family in $\overline{\mathbf{H}_{x}}$.
Proof.
Suppose that $\mathcal{F}\in\overline{\mathbf{H}_{x}}$ and $A$, $B\in\mathcal{F}$. We can see that $2\cdot f(\{A,B\},1)+2\cdot f(\{A,B\},2)\leq\overline{d(A,B)}\leq 6p+q$ so $|A\cup B|\leq\lfloor\frac{6p+q}{2}\rfloor\leq 3p+2$.
If there exists $C\in\mathcal{F}$ such that $|([n]\setminus A\setminus B)\cap C|\geq p$, then $\overline{d(\mathcal{A})}\geq\overline{d(A,B)}+2p>6p+q$ which is a contradiction. So for all $C\in\mathcal{F}$, $|([n]\setminus A\setminus B)\cap C|<p-1$ holds, therefore $|\mathcal{F}|\leq 2^{|A\cup B|}\cdot\sum_{i=0}^{p-1}{|[n]\setminus A\setminus B|\choose i}\leq 2^{3p+2}\cdot\sum_{i=0}^{p-1}{n\choose p-1}<2^{3p+2}p{n\choose p-1}$.
If $n\geq 2^{3p+2}p^{2}+p+1$ then $|\mathcal{F}|<2^{3p+2}p{n\choose p-1}\leq{n\choose p}<\sum_{i=0}^{p}{n\choose i}$, hence $\mathcal{F}$ cannot be a maximal family. ∎
Claim 4.11.
Let $x=6p+q$ and $n\geq 2^{3p+2}\cdot p^{2}+p$. If there is an $A\in\mathcal{F}\in\overline{\mathbf{H}_{x}}$ such that $|A|>p+\frac{q}{2}$ then $\mathcal{F}<\overline{h}(n,x)$.
Proof.
If $|A|\geq 3p+3$, then $\mathcal{F}\notin\overline{\mathbf{H}_{x}}$, so we can assume that $|A|\leq 3p+2$. If there exists a $B\in\mathcal{F}$ such that $|([n]\setminus A)\cap B|\geq p$, then $\overline{d(A,B)}>4p+q$, so we are done by Claim 4.10.
If there is no such set, then $|\mathcal{F}|\leq 2^{3p+2}p{n\choose p-1}<\sum_{i=0}^{p}{n\choose i}\leq\overline{h}(n,x)$. ∎
Denote the family $\mathcal{F}\cap{[n]\choose i}$ by $\mathcal{F}_{i}$.
The proof distinguishes cases according to the value of $q$:
Case 1: $q=0$ or $q=1$. By Claim 4.11, if $|\mathcal{F}|=\overline{h}(n,x)$, then all sets in $\mathcal{F}$ have at most $p$ elements. From here it is easy to see that $|\mathcal{F}|\leq\sum_{i=0}^{p}{n\choose p}$ and $\mathcal{F}=\{F:|F|\leq p\}$ is a maximal family.
Case 2: $q=2$. By Claim 4.11, if $|\mathcal{F}|=\overline{h}(n,x)$, then all sets in $\mathcal{F}$ have at most $p+1$ elements. If there are sets $A$ and $B\in\mathcal{F}$ such that $|A|=|B|=p+1$ and $|A\cap B|\leq 1$, then $\overline{d(A,B)}\geq 4p+3$, so $|\mathcal{F}|<\overline{h}(n,x)$ by Claim 4.10.
Theorem 1.6 implies $|\mathcal{F}_{p+1}|\leq{n-2\choose p+1-2}$. So $|\mathcal{F}|\leq\sum_{i=0}^{p}{n\choose p}+{n-2\choose p-1}$. The equality holds if $|F|\leq p$ or $|F|=p+1$ and $1$, $2\in F$.
Case 3: $q=4$. By Claim 4.11, if $|\mathcal{F}|=\overline{h}(n,x)$, then all sets in $\mathcal{F}$ have at most $p+2$ elements.
If there exists $A\in\mathcal{F}$ such that $|A|=p+2$, then $|A\cap B|\geq 4$ for all $B\in\mathcal{F}_{p+2}$ and $|A\cap C|\geq 2$ for all $C\in\mathcal{F}_{p+1}$ by Claim 4.10. So in this case $|\mathcal{F}_{p+1}|\leq{p+2\choose 2}{n-2\choose p-1}$ and $|\mathcal{F}_{p+2}|\leq{n-4\choose p-2}$ by Theorem 1.6. Hence we have $|\mathcal{F}|\leq{n-4\choose p-2}+{p+2\choose 2}{n-2\choose p-1}+\sum_{i=0}^{p}{n\choose p}$.
If there are no sets with $p+2$ elements, then all sets have at most $p+1$ elements. If there exist sets $A$, $B$, $C\in\mathcal{F}_{p+1}$ satisfying $d(\mathcal{A})\leq 1$, then $\overline{d(\mathcal{A})}\geq 6p+5$ by Claim 4.7. So $\mathcal{F}_{p+1}\in\mathbf{H}_{2}$. Since $n\geq 2^{3p+2}p^{2}+p+1\geq 4(p+1)^{3}$, we can use Theorem 2.5, so $\mathcal{F}$ contains at most $n-1\choose p$} sets with $p+1$ elements so $|\mathcal{F}|\leq{n-1\choose p}+\sum_{i=0}^{p}{n\choose p}$.
Comparing the two cases, we can see that the upper bound obtained in the latter case is larger, so $|\mathcal{F}|\leq{n-1\choose p}+\sum_{i=0}^{p}{n\choose p}$ and equality holds if $\mathcal{F}=\{F:|F|\leq p$ or $(|F|=p+1$ and $1\in F)\}$.
Case 4: $q=3$. Since $\mathcal{F}=\{F:|F|\leq p$ or ($|F|=p+1$ and $1\in F)\}\in\overline{\mathbf{H}_{6p+3}}$, then ${n-1\choose p}+\sum_{i=0}^{p}{n\choose p}\leq\overline{h}(n,6p+3)$.
Since $\mathcal{F}\in\overline{\mathbf{H}_{x-1}}$ implies $\mathcal{F}\in\overline{\mathbf{H}_{x}}$, then $\overline{h}(n,6p+3)\leq\overline{h}(n,6p+4)={n-1\choose p}+\sum_{i=0}^{p}{n\choose p}$.
Case 5: $q=5$. By Claim 4.11, if $|\mathcal{F}|=\overline{h}(n,x)$, then all sets in $\mathcal{F}$ have at most $p+2$ elements.
If there exist sets $A$, $B\in\mathcal{F}_{p+2}$ and a set $C\in\mathcal{F}_{p+1}$, then $|A\cap B|\geq 3$ and $|A\cap C|\geq 1$ by Claim 4.10.
There are 3 subcases we need to look at:
  1. (1)
    There exist sets $A\in\mathcal{F}_{p+2}$ and $B\in\mathcal{F}_{p+1}$ such that $|A\cap B|=1$.
  2. (2)
    There exists a set $A\in\mathcal{F}_{p+2}$ but there are no sets $B\in\mathcal{F}_{p+1}$ such that $|A\cap B|\leq 1$.
  3. (3)
    $\mathcal{F}_{p+2}$ is empty.
(1): The family $\mathcal{F}_{p+2}$ forms a $3$-intersecting family, so by Theorem 1.6, $|\mathcal{F}_{p+2}|\leq{n-3\choose p-1}$.
Since $|A\cap B|=1$, then $\overline{d(A,B)}=4p+5$, so $|C\setminus(A\cup B)|\leq p-1$ true for all $C\in\mathcal{F}_{p+1}$. From this we get that $|\mathcal{F}_{p+1}|\leq{|(A\cup B)|\choose 2}{n\choose p-1}={2p+3\choose 2}{n\choose p-1}$.
(2): $|\mathcal{F}_{p+2}|\leq{n-3\choose p-1}$ holds as before. Since $|A\cap B|\geq 2$ is true for all $B\in\mathcal{F}_{p+1}$, then $|\mathcal{F}_{p+1}|\leq{|A|\choose 2}{n\choose p-1}={p+2\choose 2}{n\choose p-1}$.
(3): If there are 3 pairwise disjoint members of $\mathcal{F}_{p+1}$, then $\overline{d(\mathcal{A})}\geq 6p+6$, so $d(\mathcal{A})\geq 1$ true for all $\mathcal{A}\subset\mathcal{F}_{p+1}$. Since $n\geq 2^{3p+2}p^{2}+p+1\geq 3(p+1)^{2}\cdot 2$, then $|\mathcal{F}_{p+1}|\leq\max\{{n\choose p+1}-{n-2\choose p+1},{(p+1)(2+1)-1\choose p+1}\}$ by using the special case of EMC [15]. If $n\geq 7p$, then $\prod_{i=n-p}^{n-2}i\geq\prod_{i=2p+4}^{3p+2}i$ and $n(n-1)-(n-p-1)(n-p-2)\geq(2p+3)(2p+2)$, so the first value is the larger and $|\mathcal{F}_{p+1}|\leq{n\choose p+1}-{n-2\choose p+1}={n-1\choose p}+{n-2\choose p}$.
Summarizing (1), (2), (3): $|\mathcal{F}|\leq\max\{{n-3\choose p-1}+{2p+3\choose 2}{n\choose p-1},{n-3\choose p-1}+{p+2\choose 2}{n\choose p-1},{n-1\choose p}+{n-2\choose p}\}+\sum_{i=0}^{p}{n\choose p}$. It can be proven by calculation that ${n-3\choose p-1}\leq{n-2\choose p}$ and ${p+2\choose 2}{n\choose p-1}\leq{2p+3\choose 2}{n\choose p-1}\leq{n-1\choose p}$ if $n\geq 8p^{3}$, so $|\mathcal{F}|\leq{n-1\choose p}+{n-2\choose p}+\sum_{i=0}^{p}{n\choose p}$. Equality holds if $\mathcal{F}=\{F:|F|\leq p$ or $(|F|=p+1$ and $|\{1,2\}\cap F|\geq 1)\}$. $\square$
Acknowledgement
The author would like to thank Gyula Katona for his helpful remarks regarding this manuscript.